OpenStudy (ttop0816):

is it right? If a line contains the point (0, -1) and has a slope of 2, then which of the following points also lies on the line? (2, 1) <--- i chose this! (1, 1) (0, 1)

OpenStudy (nirmalnema):

why you choose that

OpenStudy (ttop0816):

is it not right...? then the second one?

OpenStudy (nirmalnema):

no...means why u selected that..any reason.

OpenStudy (anonymous):

well then the line would be y=2x-1 so the only on that fits in this equation is (1,1)

OpenStudy (nirmalnema):

you can find the eq of the line by using one point form. y-y1=m(x-x1) y+1=2(x) 2x-y-1=0 now lie all the point which satisfies it that is the ans.. so the second option is corrrect. (1,1)

OpenStudy (ttop0816):

thank you!

OpenStudy (nirmalnema):

your welcome.